OpenStudy (anonymous):

Aparna walks 35km partly at the rate of 4km/hr and partly at 5km/hr. If she had walked at 5km/hr when she walked at 4km/hr and vice-versa. She would have covered 2km more in the same time. Find the time she was walking ?

OpenStudy (anonymous):

interesting question......

OpenStudy (anonymous):

I am trying to solve it from last night.

OpenStudy (anonymous):

4x + 5 y = 35 4y + 5 x = 37 x= 5 and y = 3 total = 8

OpenStudy (anonymous):

that means, she walked 8 hours

OpenStudy (anonymous):

Correct answer, but can you explain a bit.

OpenStudy (anonymous):

4 km per hour x time she walked at this rate + 5 km per hour x time she walked at this rate = total distance she walked! i.e 4x + 5y= 35

OpenStudy (anonymous):

similarly, second condition in which rates are changed.......and distance covered is two km more!

OpenStudy (anonymous):

and solving both equations we get answers

OpenStudy (anonymous):

but what partly means ?

OpenStudy (anonymous):

partly means some part of 35 kms in above situation

OpenStudy (anonymous):

nice solution saruz, I never thought of taking it as two variables.